# Infrastructure for AI Agents

### v1

@@ Abstract
Increasingly many AI systems can plan and execute interactions in open-ended environments, such as making phone calls or buying online goods. As developers grow the space of tasks that such AI agents can accomplish, we will need tools both to unlock their benefits and manage their risks. Current tools are largely insufficient because they are not designed to shape how agents interact with existing institutions (e.g., legal and economic systems) or actors (e.g., digital service providers, humans, other AI agents). For example, alignment techniques by nature do not assure counterparties that some human will be held accountable when a user instructs an agent to perform an illegal action. To fill this gap, we propose the concept of agent infrastructure: technical systems and shared protocols external to agents that are designed to mediate and influence their interactions with and impacts on their environments. Agent infrastructure comprises both new tools and reconfigurations or extensions of existing tools. For example, to facilitate accountability, protocols that tie users to agents could build upon existing systems for user authentication, such as OpenID. Just as the Internet relies on infrastructure like HTTPS, we argue that agent infrastructure will be similarly indispensable to ecosystems of agents. We identify three functions for agent infrastructure: 1) attributing actions, properties, and other information to specific agents, their users, or other actors; 2) shaping agents’ interactions; and 3) detecting and remedying harmful actions from agents. We propose infrastructure that could help achieve each function, explaining use cases, adoption, limitations, and open questions. Making progress on agent infrastructure can prepare society for the adoption of more advanced agents.
@@ 1 Introduction
Attribution (Section 3): Attributing actions, properties, and other information to agents or users Identity binding (Section 3.1): Associates an agent or its actions with a real-world identity, Such as a human or corporation. Certification (Section 3.2): Makes, verifies, and revokes claims about an agent (instance), such as what data the agent is collecting, which tools it can access, and the level of autonomy it has been authorized to exercise. Agent IDs (Section 3.3): Identifies instances of agents and links to useful attributes, such as the credentials above.
@@
As developers expand the space of tasks that AI agents can accomplish, we will need tools both to unlock their benefits and manage their risks. As an example of a potential benefit, personalized agents could help individuals make a wide variety of difficult decisions, such as what insurance to buy or which school to attend (Van Loo, 2019; Sunstein, 2024; Lazar et al., 2024). Yet, blockers such as a lack of reliability, an inability to maintain effective oversight, or the absence of recourse mechanisms could hinder adoption. Compared to the adoption of chatbots, these blockers could be more significant in the context of agents because the latter could directly cause negative consequences in the world (e.g., a mistaken financial transaction). Tools that address these blockers could enable positive uses of agents that would otherwise not materialize. On the other hand, if AI agents are adopted despite such blockers, we will need tools to prevent and respond to resulting problems. One such problem is disruptions to digital services that can result from the ability to carry out certain tasks at speed and scale, analogous to DDoS attacks. For example, an individual fradulently obtained more than 10 million dollars in royalties from Spotify by automating music creation and streaming (U.S. Attorney’s Office, 2024).
@@
The predominant focus of AI safety research is on system-level interventions, which intervene on the AI system itself to shape its behaviour. Major areas of work include goal specification and following (Hadfield-Menell et al., 2016; Christiano et al., 2017; Leike et al., 2018; Bai et al., 2022; Hua et al., 2024; Wang et al., 2024a; b; Kirk et al., 2024; Huang et al., 2024; OpenAI, 2024), adversarial robustness (Greshake et al., 2023; Tamirisa et al., 2024; Zou et al., 2024; Anil et al., 2024; Wallace et al., 2024), and cooperation (Lerer & Peysakhovich, 2019; Hu et al., 2020; Leibo et al., 2021; Dafoe et al., 2021). System-level interventions can help to improve the reliability of agents when they work and are adopted (Askell et al., 2019), but they are inapplicable to many of the problems we discuss above. For example, they by nature do not provide accountability or recourse when harm occurs. Addressing such problems will instead require structuring how agents interact with existing institutions (e.g., legal and economic systems) and actors (e.g., web service providers, humans, other AI agents).
@@
To answer this challenge, we propose the concept of agent infrastructure: technical systems and shared protocols external to agents [^2] I.e., they are not system-level interventions. that are designed to mediate and influence their interactions with and impacts on their environments [^3] See Section 7 for a comparison of agent infrastructure with adaptation interventions (Bernardi et al., 2024), which roughly include any intervention that is not a system-level intervention. These systems and protocols could be novel, or could be reconfigurations or extensions of existing systems and protocols. Examples of agent infrastructure include inter-agent communication protocols (Marro et al., 2024), IDs for agents (Chan et al., 2024b), systems for certifying an agent’s properties or behaviour, and methods for rolling back an agent’s actions (Patil et al., 2024). We include more examples in Table 1. Our notion of agent infrastructure is not about the technical systems that enable basic operation of agents (e.g., memory systems, cloud compute), although it will in practice often build upon or modify such systems. Furthermore, while our discussion will be grounded in language-model-based agents, the core ideas of agent infrastructure are largely architecture-agnostic and extend existing work across computational science, economics, and social science (Wooldridge, 2009; Perrier, 2025).
@@
AI agent safety Fine-tuning, prompting, unlearning, filtering training data Protocol to forget confidential information from an interaction, technologies to enable agent certification, distinct interaction interfaces for agents
@@
Just as the Internet relies on infrastructure like TCP (Eddy, 2022), HTTPS (Fielding et al., 2022), and BGP (Rekhter et al., 2006), we argue that agent infrastructure will likely be crucial for unlocking the benefits and managing the risks of agents. As an example of unlocking benefits, protocols that tie an agent’s actions to a user could facilitate accountability and thereby reduce blockers to agent adoption. Analogously, the ability to perform secure financial transactions over HTTPS enables an e-commerce market in the trillions of dollars (Statista, 2024). As an example of managing risks, agent infrastructure can support system-level interventions. For instance, a certification system for agents could warn actors (e.g., other agents) not to interact with agents that lack certain safeguards, just as browsers flag non-HTTPS websites. In this way, agent infrastructure can use agents’ interactions as a leverage point to improve safety: restricting an agent’s interactions correspondingly restricts the agent’s potential negative impacts.
@@ 2.1 Categories of Agent Infrastructure
Some of agent infrastructure applies to instances of agents. An agent instance is an instantiation of the underlying machine-learning model and any primitives for a particular user. Specifically, an instance would correspond to an interaction history, a set of available tools, memory, etc, much like process IDs in operating systems correspond to running instances of programs. [^5] An instance is an abstraction and may not correspond to a hardware separation. For example, inference calls for different instances could be batched together on the same piece of hardware. Agent instances are a useful concept because different instances of the same underlying model and primitives can have different users or behave differently, and therefore potentially justify different treatment.
@@ 2.2 Example
To illustrate how agent infrastructure could help unlock the benefits and manage the risks of agents, consider the task of running a company. In this hypothetical company, agents would perform or assist with the work that human employees would usually perform, such as operations, R&D, and legal tasks. Agent infrastructure could enable the following functions for this agent-run company.
@@
  • Recourse: A lack of recourse mechanisms could discourage interaction with the company’s agents. One way to address this issue is by ensuring that each agent has at least one human that can answer for the agent’s actions (Section 3.1), and making this information available to those that interact with the company’s agents or a third-party mediator.
@@ 3.1 Identity Binding
Description: AI agents are not legal entities [^7] Some argue that sufficiently capable future agents should have legal rights (Salib & Goldstein, 2024).. The ability (but not necessarily the requirement) to associate an agent’s actions with a legal entity, such as a human or a corporation, could help existing institutions accommodate agents. For example, users could be held accountable for illegal activity (e.g., fraud (Fang et al., 2024)) that they instruct their agents to carry out. Analogously, cars are associated to drivers through license plates and registries. Such binding could be helpful even when the relevant legal entity is not fully responsible for all of the agents’ actions. For example, in addition to the employee using an agent, managers reviewing the agent’s actions and the corporation hiring the employee could potentially bear some responsibility. With this motivation, we define identity binding to be a process that consists of (1) authentication of an identity and (2) linking that identity to an (instance of an) agent or its actions. Crucially, linking an identity to an agent does not mean making the identity visible publicly. For example, identity information could be accessible only to authorized parties, such as if needed for legal action, or only reveal limited information, such as the fact that a real person is responsible for the agent (Adler et al., 2024). Further below, we discuss how existing infrastructure could help accomplish (1) and (2), along with modifications that may be needed.
@@
Adoption: Firstly, users might be more likely to engage with platforms that include or require identity binding. Some evidence suggests that contractors use AI systems on platforms like Upwork (Veselovsky et al., 2023). Dharmesh (2024) is also creating a service where AI agents play the role of crowdworkers and are matched to tasks. In both of these cases, customers are likely to engage more fully if they know there is a person ultimately responsible for the service, similar to identity verification for gig economy services (Uber, 2025; DoorDash, 2025). Secondly, platforms might require identity binding to avoid Sybil-like attacks. For example, an individual recently defrauded Spotify of millions of dollars by synthetically generating music and creating multiple fake accounts to stream it (U.S. Attorney’s Office, 2024).
@@
  • How can identity binding build upon existing digital identity solutions, such as (national) digital wallets?
  • How could an authentication system be designed for agent-agent interactions and authentication of actors other than human users (e.g., for companies with digital identities (GLEIF, 2024))?
  • How can identity information be linked to an agent or its actions?
@@ 3.2 Certification
Description: Before interacting with an agent (instance), counterparties might want assurances about its operation, behaviour, or properties. For example, agents whose actions are regularly reviewed could be more trustworthy than agents who do not undergo such review. Claims about operation, behavior, or properties could come from a user, an agent deployer, or other actors (e.g., third-party auditors). Actors making such claims could include them in certificates, which could be associated with the relevant agent and made visible to counterparties interacting with it. Analagously, SSL certificates make claims to website users about domain ownership. Certification infrastructure consists of tools for making, verifying, and revoking certificates for agents. Verification includes how those viewing a certificate can verify both its claims and that it corresponds to the agent with which they are interacting.
@@
Adoption: Demand for certification is likely to vary across domains and contexts. For instance, protocols for forgetting sensitive information can enable agreements that would otherwise not be possible (Rahaman et al., 2024). Certifying that an agent can follow such protocol seems more important for high-stakes negotiations. Feasibility and economic cost could also affect demand. For example, properties that require access to interaction logs (e.g., whether a prompt injection is present) could raise privacy concerns, in the absence of privacy-preserving certification methods (Trask et al., 2023; 2024).
@@ 3.3 Agent IDs
Description: An agent ID would at minimum consist of a unique identifier [^8] A cryptographic scheme is necessary to prevent spoofing of the identifier. See Chan et al. (2024b) for more details. for an agent instance, and potentially include other information relevant to the instance, such as a system card or certifications (Chan et al., 2024b). ID-like systems exist across many domains to facilitate risk and incident management, such as serial numbers on consumer products, tail numbers on aircraft, and registration numbers for businesses. Although any AI system could have an ID, IDs could be significantly more useful for agents because they take actions in the world (see below).
@@
Counterparties that expect only to interact with agents (or which only interact with agents through specific channels [Section 4.1]) could require IDs before an interaction. Other counterparties which also interact with systems other than agents may have more difficulty doing so. For example, suppose a counterparty engages with both agents and traditional software programs. Extending a verifiable ID system from agents to all software programs would likely be infeasible, since traditional software programs do not have IDs.
@@ 4.2 Oversight Layers
Existing alternatives: Some technical frameworks already include some kind of oversight layer (Chase, 2022; Wu et al., 2023). See below for additional research questions to extend upon this work.
@@
Limitations: Oversight layers are only useful to the extent that actors use them to intervene as needed. Automation bias could reduce the ability of human actors to do so (Parasuraman & Riley, 1997; Skitka et al., 1999; Goddard et al., 2012). Weaker agents could also have difficulty in detecting problematic actions from stronger agents (Greenblatt et al., 2024).
@@
  • How can oversight layers maintain security or confidentiality (e.g., for approval of financial transaction or sending of sensitive information)?
  • Can a monitoring system accurately flag actions that are likely to be unintended? E.g., the system could draw upon past user or agent behavior.
  • Can a monitoring system accurately flag when an agent requires external intervention (e.g., additional information)?
@@
  • How can we build trusted, automated systems (Greenblatt et al., 2024; Dalrymple et al., 2024) to intervene so as to minimize the degree of human intervention required?
  • Will the market sufficiently provide for oversight layers? If not, what government action can best help incentivize their development and implementation?

### v2

@@ Abstract
AI agents plan and execute interactions in open-ended environments. For example, OpenAI’s Operator can use a web browser to do product comparisons and buy online goods. To facilitate beneficial interactions and mitigate harmful ones, much research focuses on directly modifying agent behaviour. For example, developers can train agents to follow user instructions. This focus on direct modifications is useful, but insufficient. We will also need external protocols and systems that shape how agents interact with institutions and other actors. For instance, agents will need more efficient protocols to communicate with each other and form agreements. In addition, attributing an agent’s actions to a particular human or other legal entity can help to establish trust, and also disincentivize misuse. Given this motivation, we propose the concept of agent infrastructure: technical systems and shared protocols external to agents that are designed to mediate and influence their interactions with and impacts on their environments. Just as the Internet relies on protocols like HTTPS, our work argues that agent infrastructure will be similarly indispensable to ecosystems of agents. We identify three functions for agent infrastructure: 1) attributing actions, properties, and other information to specific agents, their users, or other actors; 2) shaping agents’ interactions; and 3) detecting and remedying harmful actions from agents. We provide an incomplete catalog of research directions for such functions. For each direction, we include analysis of use cases, infrastructure adoption, relationships to existing (internet) infrastructure, limitations, and open questions. Making progress on agent infrastructure can prepare society for the adoption of more advanced agents.
@@ 1 Introduction
More general-purpose agents could automate a wide range of both beneficial and harmful tasks. Personalized agents could help individuals make a wide variety of difficult decisions, such as what insurance to buy or which school to attend (Van Loo, 2019; Sunstein, 2024; Lazar et al., 2024). Deployment of agents throughout the economy could lead to productivity growth (Korinek & Suh, 2024). Yet, barriers such as a lack of reliability, an inability to maintain effective oversight, or the absence of recourse mechanisms could hinder beneficial adoption. On the other hand, such barriers may not be critical for motivated malicious actors. Potential issues include scams (Fang et al., 2024b; Chen & Magramo, 2024) and large-scale disruption of digital services (Fang et al., 2024a; Bhatt et al., 2023; U.S. Attorney’s Office, 2024).
@@
Attribution (Section 3): Attributing actions, properties, and other information to agents or users Identity binding (Section 3.1): Associates an agent or its actions with a legal identity, such as a human or corporation. Certification (Section 3.2): Makes, verifies, and revokes claims about an agent (instance), such as what data the agent is collecting, which tools it can access, and the level of autonomy it has been authorized to exercise. Agent IDs (Section 3.3): Identifies instances of agents and links to useful attributes, such as the credentials above.
@@
To facilitate beneficial tasks and mitigate harmful ones, much AI research focuses on system-level interventions, which intervene on the AI system itself to shape its behaviour. Major areas of work include goal specification and following (Hadfield-Menell et al., 2016; Christiano et al., 2017; Leike et al., 2018; Bai et al., 2022; Hua et al., 2024; Wang et al., 2024a; b; Kirk et al., 2024; Huang et al., 2024; OpenAI, 2024), adversarial robustness (Greshake et al., 2023; Tamirisa et al., 2024; Zou et al., 2024; Anil et al., 2024; Wallace et al., 2024), and cooperation (Lerer & Peysakhovich, 2019; Hu et al., 2020; Leibo et al., 2021; Dafoe et al., 2021). System-level interventions, if adopted (Askell et al., 2019), can help to improve the reliability of agents, but they may not be sufficient for facilitating beneficial adoption or mitigating risks. For example, the difficulty of achieving adversarial robustness could mean that companies need additional assurances before adopting agents to complete economically valuable tasks. In particular, potential assurance schemes include certification of agents, insurance, or identity solutions that establish trust between different parties using agents. Such tools shape how agents interact with institutions (e.g., legal and economic systems) and other actors (e.g., web service providers, humans, other AI agents).
@@
Given the insufficiency of system-level interventions, we propose the concept of agent infrastructure: technical systems and shared protocols external to agents [^2] I.e., they are not system-level interventions. that are designed to mediate and influence their interactions with and impacts on their environments. [^3] See Section 7 for a comparison of agent infrastructure with adaptation interventions (Bernardi et al., 2024), which roughly include any intervention that is not a system-level intervention. These systems and protocols could be novel, or simply extensions of existing systems and protocols. Examples of agent infrastructure include inter-agent communication protocols (Marro et al., 2024), IDs for agents (Chan et al., 2024b), systems for certifying an agent’s properties or behaviour, and methods for rolling back an agent’s actions (Patil et al., 2024). We include more examples in Table 1. Our notion of agent infrastructure is not about the technical systems that enable basic operation of agents (e.g., memory systems, cloud compute), although it will in practice often build upon or modify such systems. Furthermore, while our discussion will be grounded in language-model-based agents, the core ideas of agent infrastructure are largely architecture-agnostic and extend existing work across computational science, economics, and social science (Wooldridge, 2009; Perrier, 2025).
@@
AI-agent safety Fine-tuning, prompting, unlearning, filtering training data Protocol to forget confidential information from an interaction, technologies to enable agent certification, distinct interaction interfaces for agents
@@
Just as the Internet relies on infrastructure like TCP (Eddy, 2022), HTTPS (Fielding et al., 2022), and BGP (Rekhter et al., 2006), we argue that agent infrastructure will likely be crucial for unlocking the benefits and managing the risks of agents. As an example of unlocking benefits, protocols that tie an agent’s actions to a user could facilitate accountability and thereby reduce barriers to agent adoption. Analogously, the ability to perform secure financial transactions over HTTPS enables an e-commerce market in the trillions of dollars (Statista, 2024). As an example of managing risks, agent infrastructure can support system-level interventions. For instance, a certification system for agents could warn actors (e.g., other agents) not to interact with agents that lack certain safeguards, just as browsers flag non-HTTPS websites. In this way, agent infrastructure can use agents’ interactions as a leverage point to improve safety: restricting an agent’s interactions correspondingly restricts the agent’s potential negative impacts.
@@ 2.1 Categories of Agent Infrastructure
Some agent infrastructure applies to instances of agents. An agent instance is an instantiation of the underlying machine-learning model and any primitives for a particular user. Specifically, an instance would correspond to an interaction history, a set of available tools, memory, etc., much like process IDs in operating systems correspond to running instances of programs. [^5] An instance is an abstraction and may not correspond to a hardware separation. For example, inference calls for different instances could be batched together on the same piece of hardware. Agent instances are a useful concept because different instances of the same underlying model and primitives can have different users or behave differently, and therefore potentially justify different treatment.
@@
Infrastructure could be implemented both within a particular organization and between organizations or actors. Consider communication protocols as an example. In the first case, a business could use communication protocols to manage its agents internally, much as infrastructure like Slack and Google docs helps human employees work together. In the second case, a common communication protocol could help agents from different organizations communicate efficiently over the internet (Marro et al., 2024; Surapaneni et al., 2025). Adoption of agent infrastructure is likely more difficult in the second case, given that different actors could use mutually incompatible communication protocols (see Section 6.1). Our presentation of agent infrastructure is meant to cover both of these cases. For each piece of infrastructure that we present, we analyze how adoption could proceed at internet-wide scale.
@@ 2.2 Example
To illustrate how agent infrastructure could work, consider the task of running a company. In this hypothetical company, agents would perform or assist with the work that human employees would usually perform, such as operations, R&D, and legal tasks. Agent infrastructure could enable the following functions for this agent-run company.
@@
  • Accountability: A lack of accountability mechanisms could discourage interaction with the company’s agents. One way to address this issue is by ensuring that each agent has at least one human that can answer for the agent’s actions (Section 3.1), and making this information available to those that interact with the company’s agents or a third-party mediator.
@@
2.3 Prioritization
We discuss some considerations for prioritization of certain types of agent infrastructure.
The effectiveness of certain infrastructure depends heavily on widespread adoption. For example, inter-agent communication protocols are only useful if both parties can use the same protocols. Agent IDs are also only useful if other parties can recognize them. On the other hand, oversight layers and rollbacks are immediately useful for addressing agent malfunction, even if few actors adopt them. For actors that are heavily resource-constrained, such considerations could suggest focusing on infrastructure that does not require widespread adoption.
Some infrastructure is only useful once agents obtain certain capabilities. For example, commitment devices are most useful when agents are capable of negotiating and carrying out complex agreements. Infrastructure for capabilities that are further off could be delayed.
Different types of infrastructure are useful for different threat models. For example, inter-agent communication protocols and commitment devices are most important for avoiding cooperation problems between agents (Hammond et al., 2025). Oversight layers and rollbacks are most useful for cases of misuse or misalignment. Other infrastructure, such as identity binding and agent IDs, can both facilitate cooperation and help to address harms.
We take no explicit stance on prioritization and leave further analysis to future work.
@@ 3.1 Identity Binding
Description: AI agents are not legal entities [^7] Some argue that sufficiently capable future agents should have legal rights (Salib & Goldstein, 2024).. The ability (but not necessarily the requirement) to associate an agent’s actions with a legal entity, such as a human or a corporation, could help existing institutions accommodate agents. For example, an agent’s VOIP calls could be tagged with the identity of the person controlling the agent, so as to enhance accountability for agent-enabled scams (Fang et al., 2024b). Analogously, cars are associated to drivers through license plates and registries. Such binding could be helpful even when the relevant legal entity is not fully responsible for all of the agents’ actions. For example, in addition to the employee using an agent, managers reviewing the agent’s actions and the corporation hiring the employee could potentially bear some responsibility. With this motivation, we define identity binding to be a process that consists of (1) authentication of an identity and (2) linking that identity to an (instance of an) agent or its actions. Crucially, linking an identity to an agent does not mean making the identity publicly visible. For example, identity information could be accessible only to authorized parties, such as if needed for legal action, or only reveal limited information, such as the fact that a real person is responsible for the agent (Adler et al., 2024). Further below, we discuss how existing infrastructure could help accomplish (1) and (2), along with modifications that may be needed.
@@
Adoption: Digital platforms are likely to be crucial for the widespread adoption of identity binding. First, users might be more likely to engage with platforms that include or require identity binding. Some evidence suggests that contractors use AI systems on platforms like Upwork (Veselovsky et al., 2023). Dharmesh (2024) is also creating a service where AI agents play the role of crowdworkers and are matched to tasks. In both of these cases, customers are likely to engage more fully if they know there is a person ultimately responsible for the service, similar to identity verification for gig economy services (Uber, 2025; DoorDash, 2025). Second, platforms might require identity binding to avoid Sybil-like attacks. For example, an individual recently defrauded Spotify of millions of dollars by synthetically generating music and creating multiple fake accounts to stream it (U.S. Attorney’s Office, 2024). Third, identity binding could be a natural extension existing single sign-on practices. For example, a Google account can already be used to log on to many non-Google-owned different websites and services.
@@
  • What kinds of actions can and should be linked to identities? For example, it is likely infeasible to link mouse clicks to agents. Would it be preferable to link groups of actions instead?
@@
  • How can identity binding build upon existing digital identity solutions, such as (national) digital wallets or single sign-on?
@@ 3.2 Certification
Description: Before interacting with an agent (instance), counterparties might want assurances about its operation, behaviour, or properties. For example, agents whose actions are regularly reviewed could be more trustworthy than agents who do not undergo such review. Claims about operation, behavior, or properties could come from a user, an agent deployer, or other actors (e.g., third-party auditors). Certificates including such claims could be associated with the relevant agent and made visible to counterparties interacting with it. Analagously, SSL certificates make claims to website users about domain ownership. Certification infrastructure consists of tools for making, verifying, and revoking certificates for agents. Verification includes how those viewing a certificate can verify both its claims and that it corresponds to the agent with which they are interacting.
@@
Adoption: Demand for certification is likely to vary across domains and contexts. For instance, protocols for forgetting sensitive information can enable agreements that would otherwise not be possible (Rahaman et al., 2024). Certifying that an agent can follow such protocol seems more important for high-stakes negotiations. Feasibility and economic cost could also affect demand. For example, properties that require access to interaction logs (e.g., whether a prompt injection is present) could raise privacy concerns, in the absence of privacy-preserving certification methods (Trask et al., 2023; 2024). That said, demand for certification could be high for agent use within government and regulated areas (e.g., health). Analogously, software products in such domains often have to meet quality standards (e.g., Health and Human Services Department & Food and Drug Administration (1996)).
@@ 3.3 Agent IDs
Description: An agent ID would at minimum consist of a unique identifier [^8] A cryptographic scheme is necessary to prevent spoofing of the identifier. See Chan et al. (2024b) for more details. for an agent instance, and potentially include other information relevant to the instance, such as a system card or certifications (Chan et al., 2024b). ID-like systems exist across many domains to facilitate identification, such as serial numbers on consumer products, tail numbers on aircraft, registration numbers for businesses, and URLs/URIs to locate resources on the internet. Although any AI system could have an ID, IDs could be significantly more useful for agents because they take actions in the world (see below).
@@
Counterparties that expect only to interact with agents (or which only interact with agents through specific channels [Section 4.1]) could require IDs before an interaction. Other counterparties which also interact with systems other than agents may have more difficulty doing so. For example, suppose a counterparty engages with both agents and traditional software programs. Extending an ID system from agents to all software programs could be infeasible, since traditional software programs do not have IDs. A potential stopgap measure could be CAPTCHA-like methods (Tatoris et al., 2025). The effectiveness of this measure depends upon whether agent behaviour is distinguishable from human and other bot behavior, such as through adversarial vulnerabilities.
@@ 4.2 Oversight Layers
Existing alternatives: Some technical frameworks for agents already include some kind of oversight layer (Chase, 2022; Wu et al., 2023). See below for additional research questions to extend upon this work.
Analogues to oversight layers are common across many settings. In banking, certain activity can trigger account freezes or requests for further review. For many digital services, logging in from an unfamiliar device triggers a request for user approval. In cyber security, endpoint detection and response tools monitor for and flag suspicious activity. Oversight layers for agents could build upon or integrate with such existing technologies.
@@
Limitations: Active attention is required if oversight layers flag potentially problematic behaviour. Humans could ignore such flags if they are too frequent or present information in a complex manner. Analogously, internet users rarely read terms of service. Humans could also have predispositions to trust automated systems without sufficient verification of their behaviour (Parasuraman & Riley, 1997; Skitka et al., 1999; Goddard et al., 2012). At the same time, AI systems could eventually be better than humans at catching problematic behaviour. For instance, language models can already help catch vulnerabilities in real-world code (Big Sleep Team, 2024). However, using more capable AI systems for monitoring incurs inference costs. Using less capable systems, or simply decreasing inference budgets, could reduce the effectiveness of monitoring. Further study is needed to understand whether the default equilibrium will be desirable.
@@
  • Who should be able to access information generated by oversight layers?
@@
  • How can we build trusted, automated systems (Greenblatt et al., 2024; Dalrymple et al., 2024) that can intervene so as to minimize the degree of human intervention required?
  • Given trade-offs between effectiveness and economic cost, will the market sufficiently provide for oversight layers? If not, what government action can best incentivize their development and implementation?